The acceleration of a particle moving along the x axis may be determined from the expression

a = btu + ctv.

What will be the dimensions of b and c in this case if

u = 10

and

v = 19?

(Use the following as necessary: L for length and T for time.)
Answer #1

Given expression is

a = btu + ctv

dimension of acceleration is

a = m/sec^2 = [L]/[T^2]

dimension of time t is

t = time = sec = [T]

Now in an expression unit of both sides should be same, So

dimension of a = dimension of btu

[L]/[T^2] = dimension of b*dimension of tu

given that u = 10

tu = t10 = [T]10

So

[L]/[T]2 = dimension of b*[T]10

[L]/[T]2*[T]10 = dimension of b

dimension of b = [L]/[T]10 + 2 = [L]/[T]12 = [L][T]-12

dimension of a = dimension of ctv

[L]/[T^2] = dimension of c*dimension of tv

given that v = 19

tv = t19 = [T]19

So

[L]/[T]2 = dimension of c*[T]19

[L]/[T]2*[T]19 = dimension of c

dimension of c = [L]/[T]19 + 2 = [L]/[T]21 = [L][T]-21

Please Upvote.